# Stability Analysis of Cosmological Perturbations in the Bumblebee Model: Parameter Constraints and Gravitational Waves

### v2

@@ III.2 Background equations
For ordinary matter, the energy density $\overline{\rho}_{m}$ is positive and corresponds to a non-negative pressure $\overline{p}_{m}$. Equation (20) shows that for a static universe ($H=0$), the energy density $\overline{\rho}_{m}$ remains constant. However, observational evidence indicates that the present universe is not only expanding but also accelerating SupernovaCosmologyProject:1998vns ; SupernovaSearchTeam:1998fmf ; Riess:1998dv . For an expanding universe, $H(t_{0})>0$ (where $t_{0}$ is the current time), which implies $\dot{\overline{\rho}}_{m}<0$. Consequently, the energy density of ordinary matter decreases with expansion, which is a natural and physically intuitive result.
@@
$\frac{1}{\kappa}\left[3H^{2}-\Lambda+3\xi b_{t}^{2}\dot{H}-\kappa(\overline{V}+2b_{t}^{2}\overline{V}_{,b^{2}})\right]=\overline{\rho}_{m},$ (23)
$\frac{1}{\kappa}\left[-3H^{2}-2\dot{H}+\Lambda+\xi b_{t}^{2}(3H^{2}+2\dot{H})+\kappa\overline{V}\right]=\overline{p}_{m},$ (24)
$b_{t}\left[-3\xi(H^{2}+\dot{H})+2\kappa\overline{V}_{,b^{2}}\right]=0,$ (25)
where $\overline{V}_{,b^{2}}\equiv\big(\partial V/\partial b^{2}\big)|_{B_{\mu}=\overline{B}_{\mu}}=\big(\partial V/\partial(B^{\mu}B_{\mu}+b^{2})\big)|_{B_{\mu}=\overline{B}_{\mu}}$, and we have used the definition of pressure $p_{m}=n\partial\rho_{m}/\partial n-\rho_{m}$. By rewriting the background equations (23)-(25), we can solve for $\xi b_{t}^{2}$, $\overline{V}_{,b^{2}}$, and $\overline{V}$, as
$\xi b_{t}^{2}$ $=$ $1+\frac{\kappa}{2\dot{H}}(\overline{p}_{m}+\overline{\rho}_{m}),$ (26)
$\overline{V}_{,b^{2}}$ $=$ $\frac{3\xi}{2\kappa}(H^{2}+\dot{H}),$ (27)
$\overline{V}$ $=$ $-\frac{\Lambda}{\kappa}+\frac{1}{\kappa}(1-\xi b_{t}^{2})(3H^{2}+2\dot{H})+\overline{p}_{m}.$ (28)
From this solution, we can see that if the Hubble parameter $H(t)$, the matter energy density $\overline{\rho}_{m}$, and the matter pressure $\overline{p}_{m}$ can be determined, the Lorentz-violating parameter $\xi b^{2}\;(=\xi b_{t}^{2})$, the vacuum potential $\overline{V}$, and $\overline{V}_{,b^{2}}$ can all be assigned definite values.
For the Hubble parameter $H$, we consider only its non-trivial solution $H=H(t)$ in this paper. The expanding universe gives $H(t_{0})>0$. The accelerating universe gives $\ddot{a}(t_{0})/a(t_{0})=H^{2}(t_{0})+\dot{H}(t_{0})>0$, which means $H^{2}(t_{0})>-\dot{H}(t_{0})$. We consider that $b_{t}$ is a non-zero constant, because the Bumblebee model is a Lorentz-violating gravity theory. Since we have not detected any Lorentz-violating effects in gravitational experiments, the Lorentz-violating parameter $\xi b^{2}$ should be very small, $|\xi b^{2}|\ll 1$. Combining Eqs. (26) and (27), we can constrain the parameter space
@@
$\frac{\overline{V}_{,b^{2}}(t_{0})}{\xi}>0,$ (30)
$\frac{\kappa}{2\dot{H}}(\overline{p}_{m}+\overline{\rho}_{m})=-1+\mathcal{O}(\xi b^{2}).$ (31)
Here, $\mathcal{O}(\xi b^{2})$ represents all higher-order quantities than $(\xi b^{2})^{0}$. According to the stability analysis of the vector perturbations in Sec. V.2, the free of tachyonic instability requires $\xi\leq 0$ (see Eq. (77)). Thus, we find $\overline{V}_{,b^{2}}(t_{0})\leq 0$. The constraint (29) is based on the premise of $\overline{\rho}_{m}\neq 0$. This constraint is consistent with both GR and cosmological observations Camarena:2019moy ; Feeney:2017sgx . Since $\dot{H}$ is seldom discussed directly in cosmology, it can be derived from the deceleration parameter $q(z)$: $\dot{H}=-(1+q)H^{2}$. The result of Ref. Camarena:2019moy gives the value of the current deceleration parameter of the universe $q_{0}=-0.55$, which indicates $\dot{H}=-0.45H^{2}$.
@@ III.3 Dark energy
In 1998, two groups, one led by Brian Schmidt and Adam Riess SupernovaSearchTeam:1998fmf and the other by Saul Perlmutter SupernovaCosmologyProject:1998vns , obtained a surprising result: the expansion rate of the universe in the past is lower than that at the present epoch, based on the sample of 16 distant and 34 nearby supernovae. Since then, dark energy has emerged as one of the central issues in theoretical physics and cosmology. The most straightforward approach to explaining dark energy is the introduction of a cosmological constant $\Lambda$ within the framework of GR, which, from the perspective of quantum field theory, is often interpreted as vacuum energy. However, there is a significant discrepancy between the cosmological constant and the prediction by quantum field theory Li:2011sd ; Cacciatori:2024ipk . In the Bumblebee model, if we attribute deviations from GR to dark energy, we can then discuss and analyze dark energy within this theoretical framework.
@@
$\overline{\rho}_{D}$ $=$ $\frac{3}{\kappa}H^{2}-\overline{\rho}_{m}=\frac{\Lambda}{\kappa}+\frac{3\xi b_{t}^{2}}{\kappa}H^{2}+\overline{V},$ (34)
$\overline{p}_{D}$ $=$ $-\frac{1}{\kappa}(3H^{2}+2\dot{H})-\overline{p}_{m}=-\frac{\Lambda}{\kappa}-\left[\frac{\xi b_{t}^{2}}{\kappa}(3H^{2}+2\dot{H})+\overline{V}\right].$ (35)
@@
$w_{D}$ $=$ $\frac{\overline{p}_{D}}{\overline{\rho}_{D}}=-1-2\xi b_{t}^{2}\frac{\dot{H}}{\kappa\overline{\rho}_{D}}.$ (36)
Since $\dot{H}\neq 0$ (see relation (29)), whether $w_{D}$ deviates from 1-1 depends on the value of $\xi b_{t}^{2}$. Typically, the term $2\xi b_{t}^{2}\dot{H}/(\kappa\overline{\rho}_{D})$ always contributes to this deviation, since the Bumblebee model is a Lorentz-violating gravity theory with $b^{2}\neq 0$ and a nonminimal coupling parameter $\xi\neq 0$. It is not difficult to see that $w_{D}$ is negative, given that $|\xi b_{t}^{2}|\ll 1$. Thus, we have $\overline{\rho}_{D}>0$, because the right side of the first equal sign of Eq. (35) implies $\overline{p}_{D}<0$. Combining the results $\dot{H}<0$ (29) and $\xi\leq 0$ (see Eq. (77)), it is easy to see that
@@
Since $\overline{\rho}_{m}>0$, $\overline{p}_{m}>0$, and $\overline{\rho}_{D}>0$, these three terms inhibit the accelerated expansion of the universe. By analyzing Eq. (35), it is easy to see that the dark energy always provides a negative pressure $\overline{p}_{D}<0$, which drives to the accelerated expansion of the universe. For the late-time cosmic acceleration, Eq. (38) gives
$\overline{p}_{D}<-\frac{1}{3}(\overline{\rho}_{D}+\overline{\rho}_{m}+3\overline{p}_{m})<0.$ (39)
If we assume that $\overline{V}=0$, then according to Eq. (35), the constraint $\xi\leq 0$ (see Eq. (77)) indicates $\Lambda>0$. It is easy to see that the cosmological constant contributes to the accelerated expansion of the universe, whereas the Bumblebee field does not.
@@ IV The tensor perturbations
$S_{t}^{(2)}=\int dtd^{3}x\frac{a^{3}}{4\kappa}\left(1-\xi b_{t}^{2}\right)\left[(\dot{h}_{+}^{2}+\dot{h}_{\times}^{2})-c_{t}^{2}\overline{g}^{zz}\big((\partial_{z}h_{+})^{2}+(\partial_{z}h_{\times})^{2}\big)\right],$ (45)
where $c_{t}^{2}=1/(1-\xi b_{t}^{2})$ is the square of the propagation speed of tensor modes. To avoid Laplacian instability, we require $c_{t}^{2}>0$, i.e.,
$1-\xi b_{t}^{2}>0.$ (46)
Obviously, this condition holds true because the Lorentz-violating parameter satisfies $|\xi b^{2}|\ll 1$. Regarding the ghost issue, since $a$ and $\kappa$ are positive, the ghost-free condition is identical to that for avoiding Laplacian instability. Therefore, the Bumblebee model is free of ghost instability in the tensor perturbations.
@@
$\ddot{h}_{w}+3H\dot{h}_{w}-\frac{\overline{g}^{zz}}{1-\xi b_{t}^{2}}\partial_{z}\partial_{z}h_{w}=0,$ (47)
where the indicator $w$ labels ++ and $\times$. Compared to the case in GR, there is only one deviation in the tensor perturbation equations of motion (47): the factor $1-\xi b_{t}^{2}$ in the third term on the left-hand side. This deviation will lead to a difference between the GW speed and the speed of light. Furthermore, the stability requirement for the vector perturbations demands that $\xi\leq 0$, resulting in the speed of tensor modes of GWs being less than 1.
On August 17, 2017, a binary neutron star coalescence candidate (GW170817) was observed through GWs by Advanced LIGO and Virgo LIGOScientific:2017vwq . About 1.7 seconds later, the Fermi Gamma-ray Burst Monitor independently detected a gamma-ray burst (GRB170817A) Savchenko:2017ffs . These observations placed a tight constraint on the speed of tensor modes of GWs LIGOScientific:2017zic ; LIGOScientific:2018dkp : $3\times 10^{-15}\leq c_{t}-1\leq 7\times 10^{-16}$. In the Bumblebee model, the Lorentz-violating parameter can be constrained using the Taylor series expansion $c_{t}=1+\xi b_{t}^{2}/2+3(\xi b_{t}^{2})^{2}/8+\mathcal{O}\big((\xi b_{t}^{2})^{3}\big)$,
$-6\times 10^{-15}\lesssim\xi b_{t}^{2}<1.4\times 10^{-15}.$ (48)
Here, the $\lesssim$” symbol and “<<” symbol both arise from the small quantity $\mathcal{O}\big((\xi b_{t}^{2})^{2}\big)$. Furthermore, the stability requirement for the vector perturbations $\xi\leq 0$ given in Eq. (77) places an additional constraint on the Lorentz-violating parameter
$-6\times 10^{-15}\lesssim\xi b_{t}^{2}\leq 0.$ (49)
@@
There are many other studies constraining the Lorentz-violating parameter in the Bumblebee model. Reference Casana:2017jkc constrains the upper bound of the Lorentz-violating parameter at the $10^{-13}$ level by time delay of light. Reference Wang:2021gtd indicates that the Lorentz-violating parameter should be negative by studying quasi-periodic oscillations frequencies in a rotating black hole. Both references consider the vacuum expectation value of the Bumblebee field to be $(0,b(r),0,0)$. For the case of a time-like vacuum expectation value $(B(t),0,0,0)$, Ref. Khodadi:2022mzt constrains the Lorentz-violating parameter by analyzing the data from Big Bang Nucleosynthesis and Gravitational Baryogenesis. This result shows that the strength of the constraint on the Lorentz-violating parameter depends on the choice of the exponent parameter $\beta$. If $0.038<\beta<0$, the Lorentz-violating parameter is positive and can be constrained at the $10^{-24}$ level. Our work differs from Ref. Khodadi:2022mzt , which employs the ansatz $B(t)\sim t^{\beta}$ for the vector field’s time evolution and introduces the term $B_{\mu}B^{\mu}R$ into the Bumblebee action. For further related studies, see Refs. Bertolami:2005bh ; Paramos:2014mda .
@@ V.1 The second-order action of the vector perturbations
$+\frac{1-\xi b_{t}^{2}}{4\kappa a}\delta^{pq}\partial_{z}\lambda_{p}\partial_{z}\lambda_{q}+\frac{\overline{J}}{2a^{2}}\overline{\rho}_{m,n}\delta^{pq}\lambda_{p}\lambda_{q}-\overline{J}\overline{\rho}_{m,n}\delta^{pq}\dot{\delta\mathcal{B}}_{p}\lambda_{q}-\frac{\xi b_{t}}{2\kappa a}\delta^{pq}\partial_{z}\lambda_{p}\partial_{z}\varsigma_{q}+\mathcal{L}_{\varepsilon}\;\;\Bigg],$
@@
$\mathcal{L}_{\varepsilon}=\frac{1-\xi b_{t}^{2}}{4\kappa}a^{3}\delta^{pq}\partial_{z}\dot{\varepsilon}_{p}\partial_{z}\dot{\varepsilon}_{q}+\frac{a}{2\kappa}\delta^{pq}\partial^{2}_{z}\left[\xi b_{t}(H\varsigma_{p}+\dot{\varsigma}_{p})-(1-\xi b_{t}^{2})(H\lambda_{p}+\dot{\lambda}_{p})\right]\varepsilon_{q}.$ (60)
@@
$-\overline{J}\overline{\rho}_{m,n}\dot{\delta\mathcal{B}}_{p}+\frac{1-\xi b_{t}^{2}}{2\kappa}a\partial_{z}^{2}\dot{\varepsilon}_{p}+\frac{\overline{J}}{a^{2}}\overline{\rho}_{m,n}\lambda_{p}-\frac{1-\xi b_{t}^{2}}{2\kappa a}\partial_{z}^{2}\lambda_{p}+\frac{\xi b_{t}}{2\kappa a}\partial_{z}^{2}\varsigma_{p}=0,$ (61)
$\partial_{z}^{2}\left(a^{2}(1-\xi b_{t}^{2})(\ddot{\varepsilon}_{p}+3H\dot{\varepsilon}_{p})+\xi b_{t}(\dot{\varsigma}_{p}+H\varsigma_{p})-(1-\xi b_{t}^{2})(\dot{\lambda}_{p}+H\lambda_{p})\right)=0,$ (62)

### v3

@@ III.2 Background equations
For ordinary matter, the energy density $\overline{\rho}_{m}$ is positive and corresponds to a non-negative pressure $\overline{p}_{m}$. Equation (20) shows that for a static universe ($H=0$), the energy density $\overline{\rho}_{m}$ remains constant. However, observational evidence indicates that the present universe is not only expanding but also accelerating Perlmutter and others (1999); Riess and others (1998, 1999). For an expanding universe, $H(t_{0})>0$ (where $t_{0}$ is the current time), which implies $\dot{\overline{\rho}}_{m}<0$. Consequently, the energy density of ordinary matter decreases with expansion, which is a natural and physically intuitive result.
@@
$\frac{1}{\kappa}\left[3H^{2}-\Lambda+3\xi b_{t}^{2}\dot{H}+3\sigma b_{t}^{2}(3H^{2}+2\dot{H})-\kappa(\overline{V}+2b_{t}^{2}\overline{V}_{,b^{2}})\right]=\overline{\rho}_{m},$ (23)
$\frac{1}{\kappa}\left[-3H^{2}-2\dot{H}+\Lambda+b_{t}^{2}(\xi+\sigma)(3H^{2}+2\dot{H})+\kappa\overline{V}\right]=\overline{p}_{m},$ (24)
$b_{t}\left[-3\xi(H^{2}+\dot{H})-6\sigma(2H^{2}+\dot{H})+2\kappa\overline{V}_{,b^{2}}\right]=0,$ (25)
where $\overline{V}_{,b^{2}}\equiv\big(\partial V/\partial b^{2}\big)|_{B_{\mu}=\overline{B}_{\mu}}=\big(\partial V/\partial(B^{\mu}B_{\mu}+b^{2})\big)|_{B_{\mu}=\overline{B}_{\mu}}$, and we have used the definition of pressure $p_{m}=n\partial\rho_{m}/\partial n-\rho_{m}$. By rewriting the background equations (23)-(25), we can obtain
$\xi b_{t}^{2}+\sigma b_{t}^{2}$ $=$ $1+\frac{\kappa}{2\dot{H}}(\overline{p}_{m}+\overline{\rho}_{m}),$ (26)
$\overline{V}_{,b^{2}}$ $=$ $\frac{3\xi}{2\kappa}(H^{2}+\dot{H})+\frac{3\sigma}{\kappa}(2H^{2}+\dot{H}),$ (27)
$\overline{V}$ $=$ $-\frac{\Lambda}{\kappa}+\frac{1}{\kappa}\left(1-(\xi+\sigma)b_{t}^{2}\right)(3H^{2}+2\dot{H})+\overline{p}_{m}.$ (28)
From these equations, we can see that if the Hubble parameter $H(t)$, the matter energy density $\overline{\rho}_{m}$, and the matter pressure $\overline{p}_{m}$ can be determined, the sum of the Lorentz-violating parameters $\xi b^{2}$ and $\sigma b^{2}$, the vacuum potential $\overline{V}$, and $\overline{V}_{,b^{2}}$ can all be assigned definite values.
For the Hubble parameter $H$, we consider only its non-trivial solution $H=H(t)$ in this paper. The expanding universe gives $H(t_{0})>0$. The accelerating universe gives $\ddot{a}(t_{0})/a(t_{0})=H^{2}(t_{0})+\dot{H}(t_{0})>0$, which means $H^{2}(t_{0})>-\dot{H}(t_{0})$. We consider that $b_{t}$ is a non-zero constant, because the Bumblebee model is a Lorentz-violating gravity theory. Since we have not detected any Lorentz-violating effects in gravitational experiments, the Lorentz-violating parameters $\xi b^{2}$ and $\sigma b^{2}$ should be very small, $|\xi b^{2}|,|\sigma b^{2}|\ll 1$. The condition $\sigma=-\tfrac{1}{2}\xi$ is ruled out by an accelerating universe, because Eq. (27) would then reduce to $H^{2}=-2\kappa\overline{V}_{,b^{2}}/(3\xi)=constant$, which is observationally excluded. Therefore, we must have $\sigma\neq-\tfrac{1}{2}\xi$. Using Eq. (26), we obtain the following constraints
@@
$\frac{\kappa}{2\dot{H}}(\overline{p}_{m}+\overline{\rho}_{m})=-1+\mathcal{O}(\xi b^{2})+\mathcal{O}(\sigma b^{2}).$ (30)
Here, $\mathcal{O}(\xi b^{2})$ and $\mathcal{O}(\sigma b^{2})$ represent all higher-order quantities than $(\xi b^{2})^{0}$ and $(\sigma b^{2})^{0}$, respectively. The constraint (29) is based on the premise of $\overline{\rho}_{m}\neq 0$. This constraint is consistent with both GR and cosmological observations Camarena and Marra (2020); Feeney et al. (2018). Since $\dot{H}$ is seldom discussed directly in cosmology, it can be derived from the deceleration parameter $q(z)$: $\dot{H}=-(1+q)H^{2}$. The result of Ref. Camarena and Marra (2020) gives the value of the current deceleration parameter of the universe $q_{0}=-0.55$, which indicates $\dot{H}=-0.45H^{2}$.
@@ III.3 Dark energy
In 1998, two groups, one led by Brian Schmidt and Adam Riess Riess and others (1998) and the other by Saul Perlmutter Perlmutter and others (1999), obtained a surprising result: the expansion rate of the universe in the past is lower than that at the present epoch, based on the sample of 16 distant and 34 nearby supernovae. Since then, dark energy has emerged as one of the central issues in theoretical physics and cosmology. The most straightforward approach to explaining dark energy is the introduction of a cosmological constant $\Lambda$ within the framework of GR, which, from the perspective of quantum field theory, is often interpreted as vacuum energy. However, there is a significant discrepancy between the cosmological constant and the prediction by quantum field theory Li et al. (2011); Cacciatori et al. (2024). In the Bumblebee model, if we attribute deviations from GR to dark energy, we can then discuss and analyze dark energy within this theoretical framework.
@@
$\overline{\rho}_{D}$ $=$ $\frac{3}{\kappa}H^{2}-\overline{\rho}_{m}=\frac{\Lambda}{\kappa}+\frac{3(\xi+\sigma)b_{t}^{2}}{\kappa}H^{2}+\overline{V},$ (33)
$\overline{p}_{D}$ $=$ $-\frac{1}{\kappa}(3H^{2}+2\dot{H})-\overline{p}_{m}=-\frac{\Lambda}{\kappa}-\left[\frac{(\xi+\sigma)b_{t}^{2}}{\kappa}(3H^{2}+2\dot{H})+\overline{V}\right].$ (34)
@@
$w_{D}$ $=$ $\frac{\overline{p}_{D}}{\overline{\rho}_{D}}=-1-2(\xi+\sigma)b_{t}^{2}\frac{\dot{H}}{\kappa\overline{\rho}_{D}}.$ (35)
Since $\dot{H}\neq 0$ (see relation (29)), whether $w_{D}$ deviates from 1-1 depends on the value of $(\xi+\sigma)b_{t}^{2}$. Typically, the term $2(\xi+\sigma)b_{t}^{2}\dot{H}/(\kappa\overline{\rho}_{D})$ always contributes to this deviation, if $b^{2}\neq 0$ and $(\xi+\sigma)\neq 0$. It is not difficult to see that $w_{D}$ is negative, given that $|(\xi+\sigma)b_{t}^{2}|\ll 1$. Thus, we have $\overline{\rho}_{D}>0$, because the right side of the first equal sign of Eq. (34) implies $\overline{p}_{D}<0$. Combining the results $\dot{H}<0$ (29) and $\xi\leq 0$ (see Eq. (76)), if $\sigma$ is zero, the equation of state of the dark energy satisfies
@@
Since $\overline{\rho}_{m}>0$, $\overline{p}_{m}>0$, and $\overline{\rho}_{D}>0$, these three terms inhibit the accelerated expansion of the universe. By analyzing Eq. (34), it is easy to see that the dark energy always provides a negative pressure $\overline{p}_{D}<0$, which drives to the accelerated expansion of the universe.
@@ IV The tensor perturbations
$S_{t}^{(2)}=\int dtd^{3}x\frac{a^{3}}{4\kappa}\left(1-(\xi+\sigma)b_{t}^{2}\right)\left[(\dot{h}_{+}^{2}+\dot{h}_{\times}^{2})-c_{t}^{2}\overline{g}^{zz}\big((\partial_{z}h_{+})^{2}+(\partial_{z}h_{\times})^{2}\big)\right],$ (43)
where $c_{t}^{2}=1+(\xi b_{t}^{2})/(1-(\xi+\sigma)b_{t}^{2})$ is the square of the propagation speed of tensor modes. Since $a$ and $\kappa$ are positive, to avoid Laplacian instability and ghost instability, we require
$\frac{\xi b_{t}^{2}}{1-(\xi+\sigma)b_{t}^{2}}>-1,$ (44)
$1-(\xi+\sigma)b_{t}^{2}>0.$ (45)
These conditions are satisfied since the Lorentz-violating parameters obey $|\xi b^{2}|,|\sigma b^{2}|\ll 1$. Therefore, the Bumblebee model is free of both Laplacian instability and ghost instability in the tensor perturbations.
@@
$\ddot{h}_{w}+3H\dot{h}_{w}-\left(1+\frac{\xi b_{t}^{2}}{1-(\xi+\sigma)b_{t}^{2}}\right)\overline{g}^{zz}\partial_{z}\partial_{z}h_{w}=0,$ (46)
where the indicator $w$ labels ++ and $\times$. Compared to the case in GR, there is only one deviation in the tensor perturbation equations of motion (46): the factor $\tfrac{\xi b_{t}^{2}}{1-(\xi+\sigma)b_{t}^{2}}$ in the third term on the left-hand side. This deviation will lead to a difference between the GW speed and the speed of light. It is straightforward to see that whether the speed of the tensor GWs deviates from the speed of light primarily depends on whether the Lorentz-violating parameter $\xi b_{t}^{2}$ is zero. Furthermore, the stability requirement for the vector perturbations demands that $\xi\leq 0$, resulting in the speed of tensor modes of GWs being less than 1.
On August 17, 2017, a binary neutron star coalescence candidate (GW170817) was observed through GWs by Advanced LIGO and Virgo Abbott and others (2017b). About 1.7 seconds later, the Fermi Gamma-ray Burst Monitor independently detected a gamma-ray burst (GRB170817A) Savchenko and others (2017). These observations placed a tight constraint on the speed of tensor modes of GWs Abbott and others (2017a, 2019): $3\times 10^{-15}\leq c_{t}-1\leq 7\times 10^{-16}$. In the Bumblebee model, the Lorentz-violating parameter can be constrained using the Taylor series expansion $c_{t}=1+\xi b_{t}^{2}/2+\mathcal{O}\big((\xi b_{t}^{2})^{n}(\sigma b_{t}^{2})^{m}\big)$, where $n+m\geq 2$,
$-6\times 10^{-15}\lesssim\xi b_{t}^{2}\lesssim 1.4\times 10^{-15}.$ (47)
Here, the $\lesssim$” symbol arises from the small quantity $\mathcal{O}\big((\xi b_{t}^{2})^{n}(\sigma b_{t}^{2})^{m}\big)$. Furthermore, the stability requirement for the vector perturbations $\xi\leq 0$ given in Eq. (76) places an additional constraint on the Lorentz-violating parameter
$-6\times 10^{-15}\lesssim\xi b_{t}^{2}\lesssim 0.$ (48)
@@
There are many other studies constraining the Lorentz-violating parameter in the Bumblebee model. Reference Casana et al. (2018) constrains the upper bound of the Lorentz-violating parameter at the $10^{-13}$ level by time delay of light. Reference Wang et al. (2022) indicates that the Lorentz-violating parameter should be negative by studying quasi-periodic oscillations frequencies in a rotating black hole. Both references consider the vacuum expectation value of the Bumblebee field to be $(0,b(r),0,0)$. For the case of a time-like vacuum expectation value $(B(t),0,0,0)$, Ref. Khodadi et al. (2023) constrains the Lorentz-violating parameter by analyzing the data from Big Bang Nucleosynthesis and Gravitational Baryogenesis. This result shows that the strength of the constraint on the Lorentz-violating parameter depends on the choice of the exponent parameter $\beta$. If $0.038<\beta<0$, the Lorentz-violating parameter is positive and can be constrained at the $10^{-24}$ level. Our work differs from Ref. Khodadi et al. (2023), which employs the ansatz $B(t)\sim t^{\beta}$ for the vector field’s time evolution. For further related studies, see Refs. Bertolami and Paramos (2005); Páramos and Guiomar (2014).
@@ V.1 The second-order action of the vector perturbations
$+\frac{1-(\xi+\sigma)b_{t}^{2}}{4\kappa a}\delta^{pq}\partial_{z}\lambda_{p}\partial_{z}\lambda_{q}+\frac{\overline{J}}{2a^{2}}\overline{\rho}_{m,n}\delta^{pq}\lambda_{p}\lambda_{q}-\overline{J}\overline{\rho}_{m,n}\delta^{pq}\dot{\delta\mathcal{B}}_{p}\lambda_{q}-\frac{\xi b_{t}}{2\kappa a}\delta^{pq}\partial_{z}\lambda_{p}\partial_{z}\varsigma_{q}+\mathcal{L}_{\varepsilon}\;\;\Bigg],$
@@
$\mathcal{L}_{\varepsilon}=\frac{1-(\xi+\sigma)b_{t}^{2}}{4\kappa}a^{3}\delta^{pq}\partial_{z}\dot{\varepsilon}_{p}\partial_{z}\dot{\varepsilon}_{q}+\frac{a}{2\kappa}\delta^{pq}\partial^{2}_{z}\left[\xi b_{t}(H\varsigma_{p}+\dot{\varsigma}_{p})-\left(1-(\xi+\sigma)b_{t}^{2}\right)\big(H\lambda_{p}+\dot{\lambda}_{p}\big)\right]\varepsilon_{q}.$ (59)
@@
$-\overline{J}\overline{\rho}_{m,n}\dot{\delta\mathcal{B}}_{p}+\frac{1-(\xi+\sigma)b_{t}^{2}}{2\kappa}a\partial_{z}^{2}\dot{\varepsilon}_{p}+\frac{\overline{J}}{a^{2}}\overline{\rho}_{m,n}\lambda_{p}-\frac{1-(\xi+\sigma)b_{t}^{2}}{2\kappa a}\partial_{z}^{2}\lambda_{p}+\frac{\xi b_{t}}{2\kappa a}\partial_{z}^{2}\varsigma_{p}=0,$ (60)
$\partial_{z}^{2}\left(a^{2}\big(1-(\xi+\sigma)b_{t}^{2}\big)(\ddot{\varepsilon}_{p}+3H\dot{\varepsilon}_{p})+\xi b_{t}(\dot{\varsigma}_{p}+H\varsigma_{p})-\big(1-(\xi+\sigma)b_{t}^{2}\big)(\dot{\lambda}_{p}+H\lambda_{p})\right)=0,$ (61)